OpenStudy (anonymous):

PLZ HELP what is the coefficient of x in the division 18x^3+12x^2-3x/6x^2

OpenStudy (anonymous):

\[\frac{18x^3+12x^2-3x}{6x^2}=\frac{18x^3}{6x^2}+\frac{12x^2}{6x^2}-\frac{3x}{6x^2}\] is a start, i.e. divide term by term

OpenStudy (anonymous):

You mean simplify?

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

it was 2 if anyone was wondering i got it wrong.

OpenStudy (anonymous):

@ShadowFang Sorry you couldn't get help, but thanks for the answer :)

OpenStudy (anonymous):

I put 2 and got it wrong, the answer is 3

OpenStudy (anonymous):

really?!? that is weird... im really sorry

OpenStudy (ckaranja):

Divide each of the terms in the numerator with \[6x^{2}\] This gives: \[\frac{18x ^{3} }{ 6x^{2} }+\frac{12x^{2} }{ 6x^{2} }-\frac{ 3x }{ 6x^{2} }\] On simplify; \[3x+2-\frac{ 3 }{ 2x }\] The co-efficient of x is 3